OpenStudy (anonymous):

Momentum problems: Please help 1) Ms. Ulrich is riding her 266 kilogram motorcycle is moving at a speed of 126m/s. what is the momentum of the cycle? 2) A river Lakes hockey player with a mass o 60kg is skating with a speed pf 12m/s. what is the player's momentum? 3)Mr.Czech is about to plow a catcher at the plate. He is running with a speed of 8m/s and has a momentum of 800 kg*m/s. What is Mr.Czech's mass? 4)if mr. Kron is the catcher and has a mass of 85kg, what will Mr. Kron's velocity be after the collision?

OpenStudy (anonymous):

All problems can be solved you momentum using momentum formula \[P = mv\] where P is momentum, m is mass of object and v is velocity of object. therefore you can plug number into equation and solve problem easily.

OpenStudy (anonymous):

thank u, would u mind showing me in question one its much easier for me to see it.

OpenStudy (anonymous):

\[P =266\left( 126m/s \right)\] what would the next step be?

OpenStudy (anonymous):

P=266*126=33516(kg*m\s)

OpenStudy (anonymous):

You need multiple the mass and speed.